(a) If P(A) = 0.48, what is the value of P(not A)?

(b) If P(not A) = 0.75, what is the value of P(not A)?

(c) Is it possible that P(A) = 17/3?

(d) Is it possible that P(A) = 3/17?

(e) Does P(A) + P(not A) always equal 1?



1
Expert's answer
2020-12-22T20:03:55-0500

(a) If P(A) = 0.48, what is the value of P(not A)?

P(not A) = 1 – P(A)

= 1 – 0.48

= 0.52

(b) If P(not A) = 0.75, what is the value of P(not A)?

P(not A) = 0.75

(c) Is it possible that P(A) = 17/3?

"P(A) = \frac{17}{3}=5.66"

No, it is not possible.

The probability of an event will not be more than 1. This is because 1 is certain that something will happen.

(d) Is it possible that P(A) = 3/17?

"P(A) = \frac{3}{17}=0.176"

Yes, it is possible. The probability of an event can be less than 1.

(e) Does P(A) + P(not A) always equal 1?

Yes, it does.